OpenStudy (anonymous):

Last 1 I promise Multiply: (2sqrt7-6sqrt6)(5sqrt7+15sqrt6)

OpenStudy (anonymous):

\[(2\sqrt{7}-6\sqrt{6}) * (5\sqrt{7} + 15\sqrt{6})\]

OpenStudy (anonymous):

-469.9999999999999

OpenStudy (anonymous):

the answer has to be in radical form

OpenStudy (anonymous):

in radical form. the answer is -470

OpenStudy (anonymous):

\[(10*7)+30\sqrt{42}-30\sqrt{42}-(90*6)\] \[70-540\] \[-470\]